# Chiral Quasicrystal Nydia Roxana Varela-Rosalesa, Michael Engela

### v1

@@ 4 Conclusions
This study demonstrated the self-assembly of a six-fold chiral quasicrystal through molecular dynamics simulations of a two-dimensional particle system. The particle, interacted via the Lennard-Jones-Gauss pair potential and were influenced by a periodic hexagonal substrate potential. Remarkably, they formed a chiral quasicrystal despite the substrate’s achiral nature. This unexpected formation underscored the complex interplay between particle interactions and external fields in quasicrystal formation causing a spontaneous breaking of mirror symmetry.
@@ Acknowledgements
Data for this article, including images/graphics and Python scripts are available at Zenodo at https://doi.org/10.5281/zenodo.13169772.

### v2

@@ 4 Conclusions
This study demonstrated the self-assembly of a six-fold chiral quasicrystal through molecular dynamics simulations of a two-dimensional particle system. The particles interacted via the Lennard-Jones-Gauss pair potential and were influenced by a periodic hexagonal substrate potential. Remarkably, they formed a chiral quasicrystal despite the substrate’s achiral nature. This unexpected formation underscored the complex interplay between particle interactions and external fields in quasicrystal formation causing a spontaneous breaking of mirror symmetry.
@@ Acknowledgements
Data for this article, including images/graphics and Python scripts are available at Zenodo at https://doi.org/10.5281/zenodo.13208339.
@@
  • Fujita 2009 N. Fujita, Acta Cryst. A, 2009, A65, 342–351.
  • Baake and Grimm 2013 M. Baake and U. Grimm, Aperiodic Order Vol. 1, sec. 6.5.1, Cambridge University Press, 2013.
  • Baake and Grimm 2017 M. Baake and U. Grimm, Aperiodic Order Vol. 2, p. 31, Cambridge University Press, 2017.